Calculate the moles of K2SO4 of a solution made by mixing 550.0mL 0.35M H2SO4 and 375.0mL 0.55M KOH. Assume 100% yield for both reactions.

                        H2SO4(aq) + 2 KOH(aq) -> 2 H2O(l) + K2SO4(aq)  

                        H2SO4(aq) + 2 H2O -> 2 H3O+(aq) + SO42-(aq)
Answer #1

Molarity(M)=number of moles of solute/Volume of solution in L

Number of moles of solute=MolarityxVolume of solution in L

1L=1000 mL

So 1mL=10-3 L

So Number of moles of solute=Molarityx(10-3 L/mL)Volume of solution in mL

So number of moles of KOH=0.55Mx(10-3 L/mL)x375mL=0.20625=0.206

Number of moles of H2SO4=0.35Mx(10-3 L/mL)x550mL=0.192

H2SO4 + KOH\rightarrowK2SO4 + H2O

So for one mole of H2SO4 reacting with 1 mole KOH, 1 mole K2SO4 is formed

Here H2SO4 is the limiting reactant as it is present in less amount, so for 0.192 moles of H2SO4, 0.192 moles KOH are consumed and 0.192 moles K2SO4 are formed.